# Smooth Tests for Normality in ANOVA

### v2

@@ 4 Data-driven choice of $K$
Such an approximation primarily accounts for the selection uncertainty of $\hat{K}$. In finite samples, under the null, the empirical probability of observing $\{\hat{K}=1\}$ may not be exactly one, as the event $\{\hat{K}=2\}$ can still occur with a small but negligible probability. Accordingly, (4.3) is obtained from $\mathrm{P}(N\hat{\Psi}_{\hat{K}}^{2}\leq x)\approx\mathrm{P}(N\hat{\Psi}_{1}^{2}\leq x,\hat{K}=1)+\mathrm{P}(N\hat{\Psi}_{2}^{2}\leq x,\hat{K}=2)$. The reliability and effectiveness of $H(x)$ are demonstrated through simulation studies in the following section.
@@ 5 Simulation
In this section, we conduct Monte Carlo experiments to examine the finite-sample performance of the proposed tests developed in Sections 3 and 4. Following Neyman, (1937) and Song and Xiao, (2022), we adopt the orthonormal Legendre polynomials on the interval $[0,1]$ for $\{\pi_{k}\}_{k=0}^{\infty}$. Under this choice, the analytical properties of the constants $c_{1k}$ and $c_{2k}$ are provided in Proposition 1 of Duchesne et al., (2016), and numerical values for $1\leq k\leq 10$ are available in Table 1 of their Supplementary Material. From a computational perspective, the test statistic $N\hat{\Psi}_{K}^{2}$, introduced in Sections 3.1 and 3.2, can be computed using a simplified expression analogous to equation (14) in Duchesne et al., (2016), which leverages the closed-form of $\bm{\Sigma}_{K}^{-1}$ (see equations (12) and (13) therein). Since Section 3.1 is a special case of Section 3.2 or 3.3, we focus our evaluation on the settings described in Sections 3.2 and 3.3 through two numerical experiments.
In the first experiment, we generate observations with different means and the same variances as follows:
$Y_{ij}^{(0,1)}\sim\mathcal{N}\left(5j,4\right),\quad Y_{ij}^{(1,1)}\sim\chi_{2}^{2}+(5j-2),\quad i=1,\ldots,jm,\quad j=1,\ldots,5,$
where $\{Y_{ij}^{(0,1)}\}_{i=1,j=1}^{jm,5}$, $\{Y_{ij}^{(1,1)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$ and $H_{1}$ respectively. Note that
@@
$\mathrm{E}[Y_{ij}^{(0,1)}]=\mathrm{E}[Y_{ij}^{(1,1)}]=5j$ and $\operatorname{Var}[Y_{ij}^{(0,1)}]=\operatorname{Var}[Y_{ij}^{(1,1)}]=4$. The sample size parameter $m$ varies from 10 to 150 in increments of 10. We set the significance level $\alpha=5\%$ and perform $500$ replications. We evaluate the performance of the test statistic $N\hat{\Psi}_{K}^{2}$ introduced in Section 3.2, with $1\leq K\leq 5$, as well as the data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$ from Section 4, where $\hat{K}$ is determined by (4.1) with $D=5$. Table 2 reports the empirical rejection rates under $H_{0}$. For fixed values of $K=1,\ldots,5$, $N\hat{\Psi}_{K}^{2}$ maintains the nominal level well, even for small sample sizes, such as $m=10$ (corresponding to a total sample size of $N=150$). For the data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$, the results indicate that the limiting $\chi^{2}_{1}$ distribution does not adequately approximate the null distribution, whereas the revised distribution $H(x)$ achieves better size control. The empirical frequency of $\hat{K}$ under $H_{0}$ and $H_{1}$ is presented in Tables 3 and 4, respectively. Under $H_{0}$, the frequency of $\{\hat{K}=1\}$ approaches 1 as $m$ increases, while under $H_{1}$, $\hat{K}$ is consistently greater than 1, which aligns with our theoretical expectations. Additionally, Figure 1 displays the sample means and corresponding error bars (sample mean $\pm$ sample standard deviation) of $\hat{K}$ across different values of $m$. Here, under our data-generating processes, under $H_{0}$, the sample mean of $\hat{K}$ remains close to 1 regardless of the sample size, and the variability diminishes as $m$ increases, indicating consistent model selection under the null. Under $H_{1}$, $\hat{K}$ increases steadily and converges to the maximum selection range $D=5$ as $m$ grows, with reduced variability. These findings align with our theoretical results and demonstrate the adaptiveness of the data-driven procedure.
@@
Table 2: Empirical rejection rates under $H_{0}$ in the first experiment, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$). The empirical power under $H_{1}$ equals 1 across all settings.
@@
Table 3: Empirical frequency of $\hat{K}$ under $H_{0}$ in the first experiment
@@
Table 4: Empirical frequency of $\hat{K}$ under $H_{1}$ in the first experiment
@@
Figure 1: The sample means and error bars of $\hat{K}$ in the first experiment
Figure 2: The sample means and error bars of $\hat{K}$ in the second experiment
In the second experiment, we generate observations with the same means and different variances as follows:
$Y_{ij}^{(0,2)}\sim\mathcal{N}\left(8,j^{2}\right),\ Y_{ij}^{(1,2)}\sim\mathcal{U}\left[8-\sqrt{3}j,8+\sqrt{3}j\right],\ i=1,\ldots,jm,\ j=1,\ldots,5,$
where $\{Y_{ij}^{(0,2)}\}_{i=1,j=1}^{jm,5}$, $\{Y_{ij}^{(1,2)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$ and $H_{1}$ respectively. Note that
$\mathrm{E}[Y_{ij}^{(0,2)}]=\mathrm{E}[Y_{ij}^{(1,2)}]=8$ and $\operatorname{Var}[Y_{ij}^{(0,2)}]=\operatorname{Var}[Y_{ij}^{(1,2)}]=j^{2}$. We adopt the same settings as in the first experiment. The results are analogous to those observed previously; see Tables 5, 6, and 7, as well as Figure 2. However, from Table 5, we observe that under $H_{1}$, the test statistic $N\hat{\Psi}_{1}^{2}$ fails to demonstrate power. This is because $\mathrm{E}[\pi_{1}(Z)]=\mathrm{E}[\pi_{1}(\Phi((Y_{ij}^{(1,2)}-8)/j))]=0$, indicating that $N\hat{\Psi}_{K}^{2}$ may lack consistency against certain directions of alternatives, particularly when $K$ is small. Correspondingly, both Table 6 and Figure 2 show that under $H_{1}$, the event $\{\hat{K}=1\}$ does not occur, which is consistent with Theorem 10 (3), where it is established that $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ under $H^{\prime}_{1}$ with $K_{0}=2$.
To summarize the simulation results, we conclude that the proposed test statistic $N\hat{\Psi}_{K}^{2}$ maintains the nominal significance level well under $H_{0}$, and demonstrates high power under the alternatives unless $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$. The data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$ using the approximated null distribution $H(x)$ performs reliably in controlling the Type  error and achieves excellent power performance.
Table 5: Empirical rejection rates in the second experiment, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
@@
Table 6: Empirical frequency of $\hat{K}$ under $H_{0}$ in the second experiment
@@
Table 7: Empirical frequency of $\hat{K}$ under $H_{1}$ in the second experiment
@@ 6 Discussion of other ANOVA models
6.1 One-way fixed effects model with different group means and variances
We first consider the one-way ANOVA model with different group means and different group variances, which is given by
@@ 6.1 One-way fixed effects model with different group means and variances
Here $\{\mu_{j}\}_{j=1}^{J},\{\sigma_{j}\}_{j=1}^{J}$ represent the potentially different group means and group standard deviations, respectively. Due to the heterogeneity across groups in model (6.1), the estimation of each group mean and group variance only depends on observations in the corresponding group, in the sense that $\hat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij},\hat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}$ for $j=1,\ldots,J$. Accordingly, we define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}_{j}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}_{j}=(Y_{ij}-\hat{\mu}_{j})/\hat{\sigma}_{j}$, and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. To implement the smooth test, one can follow the steps in Section 3 and derive a formula like regarding the estimation effects in $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij}),k=1,\ldots,K$. Then the test statistic is constructed as a quadratic form of $N^{-1/2}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$, normalized by the inverse of its asymptotic covariance matrix. It is expected that an asymptotic $\chi^{2}$ test can be obtained under mild conditions for model (6.1), which leads to a unified testing framework for one-way fixed effects models.

### v3

@@ 4 Data-driven choice of $K$
Such an approximation primarily accounts for the selection uncertainty of $\hat{K}$. In finite samples, under the null, the empirical probability of the event $\{\hat{K}=1\}$ may not be exactly $1$, as $\{\hat{K}=2\}$ can occur with a small but negligible probability. Consequently, (4.3) is derived from the approximation $\mathrm{P}(N\hat{\Psi}_{\hat{K}}^{2}\leq x)\approx\mathrm{P}(N\hat{\Psi}_{1}^{2}\leq x,\hat{K}=1)+\mathrm{P}(N\hat{\Psi}_{2}^{2}\leq x,\hat{K}=2)$. The reliability and accuracy of $H(x)$ are further illustrated through simulation studies in the next section and the appendix.
@@ 5 Simulation
In this section, we conduct Monte Carlo experiments to examine the finite-sample performance of the proposed tests developed in Sections 3 and 4. Following Neyman, (1937), Bera et al., (2013) and Song and Xiao, (2022), we adopt the orthonormal Legendre polynomials on the interval $[0,1]$ for $\{\pi_{k}\}_{k=0}^{\infty}$. Under this choice, the analytical properties of the constants $c_{1k}$ and $c_{2k}$ are provided in Proposition 1 of Duchesne et al., (2016), and numerical values for $1\leq k\leq 10$ are available in Table 1 of their Supplementary Material. From a computational perspective, the test statistic $N\hat{\Psi}_{K}^{2}$, introduced in Sections 3.1 and 3.2, can be computed using a simplified expression analogous to equation (14) in Duchesne et al., (2016), which leverages the closed form of $\bm{\Sigma}_{K}^{-1}$ [see equations (12) and (13) therein].
We first study the tests for model (3.1) through two experiments. The observations are generated as follows.
  • Experiment I:
$Y_{ij}^{(0,1)}\sim\mathcal{N}\left(5,4\right),\quad Y_{ij}^{(1,1)}\sim\chi_{2}^{2}+3,\quad i=1,\ldots,jm,\quad j=1,\ldots,5.$
  • Experiment II:
$Y_{ij}^{(0,2)}\sim\mathcal{N}\left(8,1\right),\ Y_{ij}^{(1,2)}\sim\mathcal{U}\left[8-\sqrt{3},8+\sqrt{3}\right],\ i=1,\ldots,jm,\ j=1,\ldots,5.$
Here $\{Y_{ij}^{(0,1)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(0,2)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$, while $\{Y_{ij}^{(1,1)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(1,2)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{1}$. The sample size parameter $m$ varies from 10 to 150 in increments of 10. We set the significance level $\alpha=5\%$ and perform $500$ replications. We evaluate the test statistic $N\hat{\Psi}_{K}^{2}$ from Section 3.1, with $1\leq K\leq 5$, and the data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$ from Section 4, where $\hat{K}$ is determined by (4.1) with $D=5$.
@@
Results of Experiment I are shown in Tables 2, 3, 4 and Figure 1. Table 2 reports the empirical rejection rates under $H_{0}$. For fixed $K$, $N\hat{\Psi}_{K}^{2}$ maintains the nominal level well, even for small sample sizes such as $m=10$ (corresponding to a total sample size of $N=150$). For the data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$, the limiting $\chi^{2}_{1}$ distribution underestimates the variability under $H_{0}$, while the revised distribution $H(x)$ achieves better size control. Tables 3 and 4 present the empirical frequency of $\hat{K}$ under $H_{0}$ and $H_{1}$, respectively. Under $H_{0}$, the frequency of $\{\hat{K}=1\}$ approaches $1$ as $m$ increases, whereas under $H_{1}$, $\hat{K}$ is consistently greater than $1$, in line with theoretical expectations. Figure 1 displays the sample means of $\hat{K}$ across different values of $m$, with error bars representing one standard deviation around the mean. Under $H_{0}$, the sample mean of $\hat{K}$ stays close to $1$ with diminishing variability as $m$ grows, indicating consistent model selection. Under $H_{1}$, $\hat{K}$ increases steadily and approaches the maximum $D=5$, with reduced variability, demonstrating the adaptiveness of the data-driven procedure.
The results of Experiment II are quite similar; see Tables 5, 6, and 7, as well as Figure 2. However, from Table 5, under $H_{1}$, the test statistic $N\hat{\Psi}_{1}^{2}$ fails to exhibit power because in this case, $\mathrm{E}[\pi_{1}(Z)]=\mathrm{E}[\pi_{1}(\Phi(Y_{ij}^{(1,2)}-8))]=0$. This illustrates that $N\hat{\Psi}_{K}^{2}$ may be inconsistent against certain alternatives, especially when $K$ is small. Correspondingly, both Table 6 and Figure 2 show that under $H_{1}$, $\{\hat{K}=1\}$ does not occur, consistent with Theorem 10 (3), which ensures that $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ under $H_{1}^{\prime}$ with $K_{0}=2$.
@@
Table 2: Empirical rejection rates under $H_{0}$ in Experiment I, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$). The empirical power under $H_{1}$ equals 1 across all settings.
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
10 0.050 0.048 0.048 0.044 0.040 0.056 0.034
20 0.046 0.044 0.034 0.040 0.048 0.066 0.046
30 0.042 0.036 0.044 0.068 0.068 0.060 0.046
40 0.058 0.040 0.036 0.058 0.052 0.058 0.050
50 0.064 0.044 0.054 0.054 0.062 0.068 0.060
60 0.054 0.034 0.038 0.044 0.050 0.070 0.062
70 0.056 0.038 0.048 0.046 0.052 0.056 0.048
80 0.038 0.050 0.072 0.042 0.048 0.058 0.048
90 0.062 0.066 0.052 0.064 0.058 0.048 0.040
100 0.034 0.050 0.050 0.060 0.054 0.052 0.038
110 0.054 0.032 0.054 0.060 0.050 0.038 0.036
120 0.052 0.052 0.038 0.044 0.052 0.052 0.050
130 0.052 0.050 0.064 0.060 0.058 0.068 0.064
140 0.062 0.040 0.026 0.046 0.030 0.038 0.030
150 0.054 0.044 0.062 0.034 0.058 0.056 0.048
Table 3: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment I
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
10 0.982 0.018 0 0 0
20 0.988 0.010 0 0 0.002
30 0.986 0.012 0 0.002 0
40 0.984 0.014 0.002 0 0
50 0.988 0.012 0 0 0
60 0.980 0.020 0 0 0
70 0.994 0.006 0 0 0
80 0.996 0.002 0.002 0 0
90 0.996 0.002 0.002 0 0
100 0.996 0.004 0 0 0
110 0.994 0.006 0 0 0
120 0.990 0.010 0 0 0
130 0.994 0.006 0 0 0
140 0.994 0.006 0 0 0
150 0.994 0.006 0 0 0
Table 4: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment I
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
10 0.004 0 0 0.484 0.512
20 0 0 0 0.326 0.674
30 0 0 0 0.194 0.806
40 0 0 0 0.090 0.910
50 0 0 0 0.076 0.924
60 0 0 0 0.030 0.970
70 0 0 0 0.012 0.988
80 0 0 0 0.010 0.990
90 0 0 0 0.006 0.994
100 0 0 0 0.004 0.996
110 0 0 0 0.006 0.994
120 0 0 0 0 1
130 0 0 0 0 1
140 0 0 0 0.002 0.998
150 0 0 0 0 1
Figure 1: The sample means and error bars of $\hat{K}$ in Experiment I
Table 5: Empirical rejection rates in Experiment II, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
$H_{0}$ $H_{1}$
10 0.042 0.050 0.046 0.042 0.052 0.076 0.048 0.046 0.998 0.996 0.998 0.996 0.998 0.998
20 0.034 0.042 0.026 0.050 0.060 0.070 0.054 0.018 1 1 1 1 1 1
30 0.028 0.048 0.036 0.046 0.046 0.070 0.062 0.022 1 1 1 1 1 1
40 0.054 0.054 0.046 0.050 0.060 0.062 0.052 0.042 1 1 1 1 1 1
50 0.064 0.062 0.042 0.060 0.050 0.068 0.058 0.032 1 1 1 1 1 1
60 0.032 0.054 0.060 0.040 0.040 0.078 0.064 0.026 1 1 1 1 1 1
70 0.076 0.038 0.052 0.052 0.036 0.052 0.042 0.022 1 1 1 1 1 1
80 0.040 0.046 0.034 0.048 0.058 0.074 0.068 0.028 1 1 1 1 1 1
90 0.044 0.052 0.050 0.054 0.042 0.060 0.050 0.032 1 1 1 1 1 1
100 0.038 0.068 0.050 0.046 0.046 0.064 0.056 0.028 1 1 1 1 1 1
110 0.052 0.052 0.046 0.052 0.048 0.050 0.046 0.032 1 1 1 1 1 1
120 0.052 0.042 0.038 0.034 0.052 0.040 0.036 0.030 1 1 1 1 1 1
130 0.048 0.052 0.044 0.036 0.048 0.062 0.054 0.040 1 1 1 1 1 1
140 0.062 0.058 0.056 0.038 0.038 0.046 0.042 0.044 1 1 1 1 1 1
150 0.038 0.056 0.046 0.052 0.060 0.060 0.054 0.034 1 1 1 1 1 1
Table 6: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment II
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
10 0.968 0.026 0.004 0 0.002
20 0.982 0.016 0.002 0 0
30 0.978 0.020 0.002 0 0
40 0.992 0.008 0 0 0
50 0.990 0.006 0.004 0 0
60 0.996 0.004 0 0 0
70 0.994 0.006 0 0 0
80 0.984 0.016 0 0 0
90 0.996 0.004 0 0 0
100 0.994 0.006 0 0 0
110 0.996 0.004 0 0 0
120 0.996 0.004 0 0 0
130 0.994 0.006 0 0 0
140 0.994 0.006 0 0 0
150 0.988 0.012 0 0 0
Table 7: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment II
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
10 0.002 0.984 0.012 0 0.002
20 0 0.994 0.004 0 0.002
30 0 0.990 0.008 0 0.002
40 0 0.996 0.004 0 0
50 0 0.996 0.004 0 0
60 0 0.994 0.004 0 0.002
70 0 0.998 0.002 0 0
80 0 1 0 0 0
90 0 1 0 0 0
100 0 0.996 0.004 0 0
110 0 0.998 0.002 0 0
120 0 0.994 0.006 0 0
130 0 0.998 0.002 0 0
140 0 0.998 0.002 0 0
150 0 0.996 0.004 0 0
Figure 2: The sample means and error bars of $\hat{K}$ in Experiment II
The next two experiments are designed to evaluate the tests for models (3.9) and (3.10), respectively. All settings are inherited from the previous two experiments, except that the data-generating process is modified accordingly.
  • Experiment III:
$Y_{ij}^{(0,3)}\sim\mathcal{N}\left(5j,4\right),\quad Y_{ij}^{(1,3)}\sim\chi_{2}^{2}+(5j-2),\quad i=1,\ldots,jm,\quad j=1,\ldots,5.$
  • Experiment IV:
$Y_{ij}^{(0,4)}\sim\mathcal{N}\left(8,j^{2}\right),\ Y_{ij}^{(1,4)}\sim\mathcal{U}\left[8-\sqrt{3}j,8+\sqrt{3}j\right],\ i=1,\ldots,jm,\ j=1,\ldots,5.$
Here $\{Y_{ij}^{(0,3)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(0,4)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$, while $\{Y_{ij}^{(1,3)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(1,4)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{1}$. Note that in Experiment III, $\mathrm{E}[Y_{ij}^{(0,3)}]=\mathrm{E}[Y_{ij}^{(1,3)}]=5j$ and $\operatorname{Var}[Y_{ij}^{(0,3)}]=\operatorname{Var}[Y_{ij}^{(1,3)}]=4$, which corresponds to model (3.9); in Experiment IV, $\mathrm{E}[Y_{ij}^{(0,4)}]=\mathrm{E}[Y_{ij}^{(1,4)}]=8$ and $\operatorname{Var}[Y_{ij}^{(0,4)}]=\operatorname{Var}[Y_{ij}^{(1,4)}]=j^{2}$, which corresponds to model (3.10). The conclusions from Experiments III and IV are analogous to those from Experiments I and II, respectively (see Tables 813 and Figures 34), thus validating our unified testing framework. To examine the robustness of our findings with respect to the number of groups, we additionally consider a variant of Experiment III with $J=10$ (referred to as Experiment III). The results corroborate the main conclusions, as reported in A.
Table 8: Empirical rejection rates under $H_{0}$ in Experiment III, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$). The empirical power under $H_{1}$ equals 1 across all settings.
@@
Table 9: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment III
@@
Table 10: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment III
@@
Figure 3: The sample means and error bars of $\hat{K}$ in Experiment III
Figure 4: The sample means and error bars of $\hat{K}$ in Experiment IV
In summary, the experiments demonstrate that the proposed tests in Section 3 maintain the nominal significance level well under the null, and exhibit high power under the alternatives, except when $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$. The data-driven tests with the approximated null distribution $H(x)$ reliably control the Type I error and achieve excellent power performance.
Table 11: Empirical rejection rates in Experiment IV, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
@@
Table 12: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment IV
@@
Table 13: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment IV
@@ 6 Discussion of other ANOVA models
6.1 One-way fixed effects model with heterogeneous means and variances
We first consider the one-way ANOVA model with potentially different group means and variances, given by
@@ 6.1 One-way fixed effects model with heterogeneous means and variances
Here $\{\mu_{j}\}_{j=1}^{J}$ and $\{\sigma_{j}\}_{j=1}^{J}$ denote the group-specific means and standard deviations. Owing to heterogeneity across groups, the estimators of each group mean and variance rely solely on the corresponding observations, in the sense that $\hat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij},\hat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}$ for $j=1,\ldots,J$. Accordingly, we define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}_{j}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}_{j}=(Y_{ij}-\hat{\mu}_{j})/\hat{\sigma}_{j}$, and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. To implement the smooth test, one may follow the procedure in Section 3 and derive an expression for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$, $k=1,\ldots,K$, that accounts for the estimation effects. The test statistic is then constructed as a quadratic form of $N^{-1/2}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$, normalized by the inverse of its asymptotic covariance matrix. Under mild conditions, an asymptotic $\chi^{2}$ test can be established for model (6.1), thereby extending the unified framework to this case.